ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 103, 1998

Ambartzumian, R. V. and V. K. Oganian Parametric versions of Hilbert’s fourth problem

41

Assani, I. Multiple recurrence and almost sure convergence for weakly mixing dynamical systems

111

Berele, A. and A. Regev Codimensions of products and of intersections of verbally prime T -ideals

17

Björner, A. and K. S. Sarkaria The zeta function of a simplicial complex

29

Cartwright, D. I. and T. Steger A family of Ã_{n}-groups  125

Casazza, P. G. and N. J. Kalton Uniqueness of unconditional bases in Banach spaces

141

Chan, H. H. and M. L. Lang Ramanujan’s modular equations and Atkin–Lehner involutions

1

Dèbes, P. and J-C. Douai Local-global principles for algebraic covers

237

Douai, J-C. See P. Dèbes

du Sautoy, M. P. F. and G. R. Everest Integrating on p-adic Lie groups

207

Everest, G. R. See M. P. F. du Sautoy

Jia, R-Q. Shift-invariant spaces and linear operator equations  259

Kalton, N. J. See P. G. Casazza

Lang, M. L. See H. H. Chan

Lichtman, A. I. The residual nilpotence of verbal wreath products  319
Madden, K. M. The giant spiral  67

Oganian, V. K. See R. V. Ambartzumian

Regev, A. See A. Berele

Sarkaria, K. S. See A. Björner

Shvartsman, P. A geometrical approach to the K-divisibility problem

289

Steger, T. See D. I. Cartwright

Thomas, S. The automorphism tower problem II  93

Wehrung, F. Non-measurability properties of interpolation vector spaces

177

k